40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Sep 22 13:56:56 2011, kiwamu, Update, ASC, C1ASS : status update 
    Reply  Thu Sep 22 17:48:20 2011, kiwamu, Update, SUS, bad oplev on ETMY OLETMY.png
       Reply  Thu Sep 22 20:12:54 2011, kiwamu, Update, SUS, ETMY oplev whitening engaged OLETMY_WF.png
    Reply  Mon Sep 26 12:41:27 2011, kiwamu, Update, ASC, C1ASS : status update 
       Reply  Fri Sep 30 01:18:39 2011, kiwamu, Update, ASC, C1ASS : status update 
Message ID: 5521     Entry time: Thu Sep 22 17:48:20 2011     In reply to: 5518     Reply to this: 5523
Author: kiwamu 
Type: Update 
Category: SUS 
Subject: bad oplev on ETMY 

It turned out the oplev controls on ETMY were just bad.

It looks like the whitening filters have been OFF and because of that the resultant open-loop was not crossing the unity gain.

I will check the whitening filters.

 

(open-loop transfer function)

The blue dots are the measured data points and the green curve is the fit.

Apparently the open-loop doesn't go above the unity gain, so the oplev had been doing nothing.

If we try to increase the overall gain it will oscillate because of the phase delay of more than 180 deg around 3 Hz.

The red curve is the expected one with the whitening filters (WFs) properly engaged.

Note that WF are supposed to have two zeros at 1 Hz and two poles at 10 Hz.

 OLETMY.png

Quote from #5518
(to do)
 + optimization of the ETMY oplevs and OSEM damping.

ELOG V3.1.3-